# Exploring a Patch-Wise Approach for Privacy-Preserving Fake ID Detection

### v1

@@ 6 Conclusion
This paper has presented a novel patch-wise approach for privacy-preserving fake ID detection, exploring several configurations in terms of performance and privacy. Due to the lack of public databases that contain both real and fake ID documents, we have acquired a novel databases comprising real ID documents from 30 subjects in total. In addition, fake ID documents using print and screen methods are included in the public database.
Through an in-depth experimental framework, we have validated our proposed method considering intra- and cross-database scenarios. In particular, our optimal configuration is based on DINOv2 model, with a configuration based on $64\times 64$ patches and fully-anonymized ID. With this setup, and over a different database not seen in training (DLC-2021), our proposed method has been able to achieve EER values of 13.91% and 0% for the analysis at patch level and the whole ID document level, respectively, proving to generalize well to other PAIs and conditions. However, we are aware that these good results might be produced due to the lack of public databases as they do not cover all possible real-life scenarios. Future work will be oriented to increase the size and variability of our public database.
@@ Acknowledgements
This study is supported by INTER-ACTION (PID2021-126521OBI00 MICINN/FEDER), Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), and PowerAI+ (SI4/PJI/2024-00062, funded by Comunidad de Madrid, Spain through the grant agreement for the promotion of research and technology transfer at UAM).

### v2

@@ 6 Conclusion
We presented FakeIDet, a novel patch-wise approach for privacy-preserving fake ID detection and explored several configurations in terms of performance and privacy. Due to the lack of public databases that contain both real and fake IDs, we acquired a new database (FakeIDet-db), which includes real IDs from 30 subjects plus fake IDs using print and screen methods.
Through an in-depth experimental framework, we have validated our proposed method considering intra- and cross-database scenarios. In particular, our optimal configuration is based on the DINOv2 model, with a configuration based on $\times 64$ 64 patches and fully-anonymized ID. With this setup and over a different database not seen in training (DLC-2021), our proposed method has been able to achieve EER values of 13.91% and 0% for the analysis at patch level and the whole ID level, respectively, proving to generalize well to other PAIs and conditions. However, we are aware that these good results might be produced due to the lack of public databases, as they do not cover all possible real-life scenarios. Future work will be oriented towards increasing the size and variability of our public database.
@@ Acknowledgements
Support by INTER-ACTION (PID2021-126521OBI00 MICINN/FEDER), M2RAI (PID2024-160053OB-I00 MICIU/FEDER), Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), FakeIDet (FUAM-Veridas R&D Agreement signed July 2025), and PowerAI+ (SI4/PJI/2024-00062 Comunidad de Madrid and UAM).